# Seek in the Dark: Reasoning via Test-Time Instance-Level Policy Gradient in Latent Space

### v1

@@ Appendix D Qualitative Analysis and Case Studies
To gain a deeper understanding of model-generated responses, we analyze the distribution of the first three tokens produced under Prompt 1 in the GSM8K dataset Cobbe et al. (2021), using the Llama3.1-8B-Instruct model. These distributions are visualized in the word cloud shown in Figure 4. The appearance of nonsensical or obscure tokens—such as “thecy,” “theella,” and unrelated proper names like “Mark” and “John”—suggests that the reasoning processes employed by language models can diverge significantly from human reasoning, despite generating superficially plausible outputs. Notably, consistent patterns emerge: the first generated tokens are often prepositions (e.g., “let”), the second tokens are primarily verbs (e.g., “find,” “solve”), and the third tokens are predominantly proper names, These findings suggest that optimal reasoning paths for language models may differ from human strategies and could be more effectively modeled in latent space.
We present several representative cases in Tables 19, 20, 21, 22, 23, 24 and 25 to demonstrate the divergence between machine reasoning patterns and human cognition processes. These cases reveal that while our method (OURS) generates linguistically anomalous expressions such as "total downloads of downloads", "let need calculate to calculate" and "let’ll more understand it", it still achieves correct solutions. This suggests that even if the decoded tokens from our updated embeddings lack syntactic and semantic coherence, the logic and answer of the final generated output remain accurate. Notably, although the baseline model produces erroneous answers that appear fundamentally distinct from ours (OURS) in textual form, their latent representations exhibit structural proximity - through minimal adjustments to embeddings in the latent space, the original output (ORIGINAL MODEL) can be transfromed into correct solutions (OURS). These empirical evidences yield two critical implications. First, it suggests that the reasoning paths employed by machines are fundamentally distinct from those of humans. Second, while model errors may appear substantially deviant from correct answers in textual outputs, they may maintain significant proximity to valid solutions in latent space.
@@
Appendix E Performance vs Fraction Ratio
Figure Figure 5 illustrates the performance trends of various instruction-tuned language models across different fraction ratios, which range from 0.1 to 0.8. Performance is reported as a percentage and plotted on the y-axis, while the fraction ratio is shown on the x-axis.
Figure 5: Performance vs Fraction Ratio
Appendix F Limitations of Outcome Reward Model
The reward function of using a outcome reward model is defined as follows:
$\tilde{x}\sim\pi(\tilde{x}|x)$ (13)
$R(x,c)\sim\text{RewardModel}(\cdot|\tilde{x},c)$ (14)
The performance is described in Liu et al. (2024b), with results presented in Table 26. Although this model achieves an average score of 60.16—representing an improvement of 1.91 points over the Chain-of-Thought (CoT) method—it remains significantly inferior to the performance attained using a self-rewarding mechanism, which reaches 71.74. These results highlight the limitations of current outcome-based reward models in scenarios lacking ground-truth data, underscoring the need for the development of more robust and effective outcome reward models.
Methods Model Qwen2 Qwen2.5 LLaMA3.1 Mistral Avg
7B 1.5B 7B 14B 8B 7B
GSM8k: Prompt 2
CoT 65.20 15.31 66.41 91.81 69.07 41.70 58.25
LatentSeek (Math) 61.18 28.28 66.64 91.05 76.80 37.00 60.16
LatentSeek (Self) 80.21 44.20 85.06 92.72 83.70 44.58 71.74
Appendix G Computational Resources
@@
Appendix H Ethics and Societal Impact
This work focuses on improving the reasoning capabilities of large language models (LLMs). The research is methodological in nature and does not involve human subjects, sensitive data, or deployment in real-world applications. Our contributions are confined to improving core algorithmic aspects of LLM reasoning and do not introduce new data that could raise concerns regarding privacy, bias, or misuse. While we recognize that LLMs can have broader societal impacts, particularly when used in downstream applications, our work does not directly engage with these deployment scenarios. We also note that enhanced reasoning capabilities may indirectly influence downstream model behavior. However, the improvements described in this paper are academic-purpose and do not facilitate manipulation, deception, or unethical use of LLMs. Overall, we believe that our research poses no direct ethical or societal risks and is aligned with the responsible development of trustworthy AI systems.
Appendix I Limitations and future works
Reward Models.
One limitation of our current approach lies in the use of a self-rewarding mechanism. Ideally, employing a more principled outcome-based reward model would be preferable. However, our experiments indicate that existing outcome reward models are not yet sufficiently effective, suggesting a promising direction for future research.
Latent Optimization.
We adopt basic policy gradient methods in our implementation, deliberately leaving the exploration of more advanced reinforcement learning algorithms—such as Proximal Policy Optimization (PPO)—to future work by the research community.
Large Base Model.
Our experiments are conducted using 14B-parameter models, constrained by available computational resources. Scaling the approach to larger base models remains an important avenue for future investigation.

### v2

@@ Appendix H Qualitative Analysis and Case Studies
To gain a deeper understanding of model-generated responses, we analyze the distribution of the first three tokens produced under Prompt 1 in the GSM8K dataset Cobbe et al. (2021), using the Llama3.1-8B-Instruct model. These distributions are visualized in the word cloud shown in Figure˜7. The appearance of nonsensical or obscure tokens—such as “thecy,” “theella,” and unrelated proper names like “Mark” and “John”—suggests that the reasoning processes employed by language models can diverge significantly from human reasoning, despite generating superficially plausible outputs. Notably, consistent patterns emerge: the first generated tokens are often prepositions (e.g., “let”), the second tokens are primarily verbs (e.g., “find,” “solve”), and the third tokens are predominantly proper names, These findings suggest that optimal reasoning paths for language models may differ from human strategies and could be more effectively modeled in latent space.
We present several representative cases in Tables˜23, 24, 25, 26, 27, 28 and 29 to demonstrate the divergence between machine reasoning patterns and human cognition processes. These cases reveal that while our method (OURS) generates linguistically anomalous expressions such as "total downloads of downloads", "let need calculate to calculate" and "let’ll more understand it", it still achieves correct solutions. This suggests that even if the decoded tokens from our updated embeddings lack syntactic and semantic coherence, the logic and answer of the final generated output remain accurate. Notably, although the baseline model produces erroneous answers that appear fundamentally distinct from ours (OURS) in textual form, their latent representations exhibit structural proximity - through minimal adjustments to embeddings in the latent space, the original output (ORIGINAL MODEL) can be transfromed into correct solutions (OURS). These empirical evidences yield two critical implications. First, it suggests that the reasoning paths employed by machines are fundamentally distinct from those of humans. Second, while model errors may appear substantially deviant from correct answers in textual outputs, they may maintain significant proximity to valid solutions in latent space.
@@
Appendix I Computational Resources
@@
Appendix J The Use of Large Language Models (LLMs)
Our use of Large Language Models (LLMs) was strictly limited to polishing the language and generating figures for the manuscript. All underlying research and intellectual content of this paper, including the LatentSeek framework, its theoretical foundations, experimental design, and the analysis of results, was completed entirely by the authors without assistance from LLMs.